OpenStudy (anonymous):

Solve. 64^x = 16^x–1

OpenStudy (anonymous):

Find a common base which is 4 4^4x=4^2x-2, and now bring the exp down and solve 4x=2x-2 2x=-2 x=-1

OpenStudy (anonymous):

If you need me to explain what I did, I can do that.

OpenStudy (anonymous):

the -1 is part of the exponent?, the problem gets significantly difficult if it is not

OpenStudy (anonymous):

yes it is i should have put a parenthesis thank you so much

OpenStudy (anonymous):

lozl taco bell man thank you very much :)